# Discrete conformal structures on surfaces with boundary (I)—Classification

### v3

@@ Abstract.
In this paper, we introduce the discrete conformal structures on surfaces with boundary in an axiomatic approach, which ensures that the Poincaré dual of an ideally triangulated surface with boundary has a good geometric structure. Then we classify the discrete conformal structures on surfaces with boundary, which turns out to unify and generalize Guo-Luo’s generalized circle packings [8], Guo’s vertex scalings [7] and Xu’s partial discrete conformal structures [14] on surfaces with boundary. Motivated by [2, 18], we further study the relationships between the discrete conformal structures on surfaces with boundary and the hyperbolic trigonometry. Unexpectedly, we find that some subclasses of the discrete conformal structures on surfaces with boundary are closely related to the twisted generalized hyperbolic triangles introduced by Roger-Yang [11], which does not appear in the case of closed surfaces. Finally, we study the relationships between the discrete conformal structures on surfaces with boundary and the 3-dimensional hyperbolic geometry by constructing ten types of generalized hyperbolic tetrahedra. Some new generalized hyperbolic tetrahedra recently introduced by Belletti-Yang [1] naturally appears in the constructions.
@@ 1.2. Classification of discrete conformal structures on surfaces with boundary
Remark 1.7.
Note that we require the condition (1) holds for any right-angled hyperbolic hexagon $\{ijk\}\in F$ in $(\Sigma,\mathcal{T})$. By Theorem 1.4, the discrete conformal structures (4), (5) and (6) exists in $(\Sigma,\mathcal{T})$ only if $\frac{\sinh d_{ij}}{\sinh d_{ji}}>0$ for any right-angled hyperbolic hexagon $\{ijk\}$. Furthermore, if $d>0$ for any right-angled hyperbolic hexagon $\{ijk\}$, then the condition (1) holds. Thus it is possible that there exist only the discrete conformal structures (4), (5) or (6) on an ideally triangulated surface with boundary. However, by the proof of Theorem 1.4, the discrete conformal structures (4), (5) and (6) can not exist simultaneously on the same ideally triangulated surface with boundary. Please refer to Subsection 2.4 for more details on this. According to Theorem 1.4 again, the discrete conformal structures (7), (8) and (9) exist in $(\Sigma,\mathcal{T})$ only if $\frac{\sinh d_{ij}}{\sinh d_{ji}}<0$ for any right-angled hyperbolic hexagon $\{ijk\}$. This implies that the discrete conformal structures (7), (8) and (9) can not exist alone on an ideally triangulated surface with boundary. Otherwise, we have $\frac{\sinh d_{ij}}{\sinh d_{ji}}<0$$\frac{\sinh d_{ki}}{\sinh d_{ik}}<0$ and $\frac{\sinh d_{jk}}{\sinh d_{kj}}<0$, which implies $\frac{\sinh d_{ij}\sinh d_{jk}\sinh d_{ki}}{\sinh d_{ji}\sinh d_{ik}\sinh d_{kj}}<0$. This contradicts the condition (1). Therefore, we have to consider the mixed type of discrete conformal structures, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (4) and (7), (5) and (8), or (6) and (9), such that the condition (1) holds. One can check that the discrete conformal structures (4) and (8), (4) and (9), (5) and (7), (5) and (9), (6) and (7), and (6) and (8) can not exist simultaneously on the same ideally triangulated surface with boundary $(\Sigma,\mathcal{T})$ by the proof of Theorem 1.4.
@@ Acknowledgements
This work was once presented on the workshop “Discrete Conformal Geometry” taking place at Wuhan University in China from June 21st to June 24th, 2024. During the workshop, Professor Feng Luo at Rutgers University and Professor Tian Yang at Texas A&M University suggested the relationships between the twisted generalized triangles and the discrete conformal structures on surfaces with boundary. The authors thank Professor Feng Luo and Professor Tian Yang for their suggestions.
@@ 2.1. Lorentzian cross product
For any two vectors $x,y\in\mathbb{H}^{2}$, there is a unique geodesic through $x$ and $y$. The arc length of the geodesic between $x$ and $y$ is defined to be the hyperbolic distance between $x$ and $y$, denoted by $d_{\mathbb{H}}(x,y)$. We use $\mathrm{Span}(x)$ and $\mathrm{Span}(x,y)$ to denote the vector subspace spanned $x$ and $x,y$, respectively. If $x\in\mathbb{R}^{3}$ is a space-like vector, then the Lorentzian complement $x^{\bot}:=\{w\in\mathbb{R}^{3}|x*w=0\}$ of $\mathrm{Span}(x)$ is a 2-dimension vector subspace of $\mathbb{R}^{3}$ and $x^{\bot}\cap\mathbb{H}^{2}$ is non-empty. And the hyperbolic distance between $x$ and $y\in\mathbb{H}^{2}$ is defined to be $d_{\mathbb{H}}(y,x^{\bot})=\inf\{d_{\mathbb{H}}(y,z)|z\in x^{\bot}\cap\mathbb{H}^{2}\}$. Let $x,y\in\mathbb{R}^{3}$ be space-like vectors. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is non-empty, then the hyperbolic distance between $x$ and $y$ is defined by $d_{\mathbb{H}}(x^{\bot},y^{\bot})=\inf\{d_{\mathbb{H}}(z,w)|z\in x^{\bot}\cap\mathbb{H}^{2},w\in y^{\bot}\cap\mathbb{H}^{2}\}$. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is empty, then $x^{\bot}$ and $y^{\bot}$ intersect in an angle $\angle(x^{\bot},y^{\bot})$ in $\mathbb{H}^{2}$.
@@ 2.3. Conformal variations of a right-angled hyperbolic hexagon
For a right-angled hyperbolic hexagon $\{ijk\}\in F$, there exists a unique hyper-ideal hyperbolic triangle $\{ijk\}$ corresponding to it with the vertices corresponding to the boundary arcs. We consider the effect of perturbing $f_{k}$ conformally while keeping $f_{i}$ and $f_{j}$ fixed, i.e., $\delta f_{i}=\delta f_{j}=0$. This is called the $f_{k}-$conformal variation by Glickenstein-Thomas [6]. In [6], Glickenstein-Thomas added a condition to a hyperbolic triangle in a closed surface, requiring that the vertex $v_{k}$ and face center $c_{ijk}$ of a hyperbolic triangle should be in a line including the vertex $\overline{v}_{k}$ of the new hyperbolic triangle under the $f_{k}-$conformal variation. We change the condition “in a line” in Glickenstein-Thomas [6] to the condition “in a 2-dimensional vector subspace of $\mathbb{R}^{3}$” for a hyper-ideal hyperbolic triangle in an ideally triangulated surface with boundary. In the Klein model, this is almost equivalent to the condition introduced by Glickenstein-Thomas, except that the vertices $v_{k}$ and $\overline{v}_{k}$ are spacelike. Please refer to Figure 2. Note that if the vertex $v_{i}$ of a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ is fixed, then the position of boundary component $i=v_{i}^{\bot}\cap\mathbb{H}^{2}$ will not move, while the arc length may change.
@@
$\log\frac{\sinh^{2}d_{ij}}{\sinh^{2}d_{ji}}+\log\frac{\sinh^{2}d_{jk}}{\sinh^{2}d_{kj}}+\log\frac{\sinh^{2}d_{ki}}{\sinh^{2}d_{ik}}=0.$
Then $c_{ij}+c_{jk}+c_{ki}=0$ by (30). Hence, $C_{ij}+C_{jk}+C_{ki}=0$ and $C_{rs}+C_{sr}=0$ for $\{r,s\}\subseteq\{i,j,k\}$. This gives the discrete conformal structure (6).
@@
This gives the discrete conformal structure (4).
@@
This gives the discrete conformal structure (7).
@@ 3. Relationships with 3-dimensional hyperbolic geometry
Denote the hyperbolic line dual to $O,v_{i},v_{j}$ by $L_{O},L_{i},L_{j}$ respectively. According to Remark 1.7, the discrete conformal structures (4), (5) and (6) can exist alone on an ideally triangulated surface with boundary. Hence, the three lateral generalized triangles with the cosine law from (4), (5) or (6) of a hyperbolic tetrahedron are the same.
@@
Now we give the geometric explanations of the discrete conformal structures (7), (8) and (9). According to Remark 1.7 again, the discrete conformal structures (7), (8) and (9) can not exist alone on an ideally triangulated surface with boundary. And we consider the mixed type of discrete conformal structures, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (4) and (7), (5) and (8) or (6) and (9). To ensure the condition (1) holds, there must be one of $l_{ij},l_{jk},l_{ki}$ from (4) and the other two from (7), one from (5) and the other two from (8), or one from (6) and the other two from (9). In other words, in a hyperbolic tetrahedron, there exist one lateral generalized triangle with the cosine law from (4) and the other two from (7), one lateral generalized triangle with the cosine law from (5) and the other two from (8), or one lateral generalized triangle with the cosine law from (6) and the other two from (9). For the discrete conformal structures (4), (5) and (6), the lateral generalized triangles are still shown in Figure 4. However, for the discrete conformal structures (7), (8) and (9), the lateral generalized triangles are some twisted hyperbolic triangles introduced by Roger-Yang [11]. Please refer to Figure 5. One can also refer to [1, 11] for more information on twisted hyperbolic triangles.
@@
Figure 6. Schematic diagram of fifth generalized hyperbolic tetrahedra and the corresponding lateral generalized triangles in the Kelin model.

### v4

@@ Abstract.
In this paper, we introduce the discrete conformal structures on surfaces with boundary in an axiomatic approach, which ensures that the Poincaré dual of an ideally triangulated surface with boundary has a good geometric structure. Then we classify the discrete conformal structures on surfaces with boundary, which turns out to unify and generalize Guo-Luo’s generalized circle packings [8], Guo’s vertex scalings [7] and Xu’s partial discrete conformal structures [14] on surfaces with boundary. This generalizes the results of Glickenstein-Thomas [6] on closed surfaces to surfaces with boundary. Motivated by [2, 18], we further study the relationships between the discrete conformal structures on surfaces with boundary and the hyperbolic trigonometry. Unexpectedly, we find that some subclasses of the discrete conformal structures on surfaces with boundary are closely related to the twisted generalized hyperbolic triangles introduced by Roger-Yang [11], which does not appear in the case of closed surfaces. Finally, we study the relationships between the discrete conformal structures on surfaces with boundary and the 3-dimensional hyperbolic geometry by constructing ten types of generalized hyperbolic tetrahedra. Some new generalized hyperbolic tetrahedra recently introduced by Belletti-Yang [1] naturally appears in the constructions.
@@ 1.2. Classification of discrete conformal structures on surfaces with boundary
Furthermore, we have
(a):
any one of the discrete conformal structures (4), (5) or (6) can exist alone on an ideally triangulated surface with boundary, while the discrete conformal structures (4), (5) and (6) can not exist simultaneously in pairs on the same ideally triangulated surface with boundary.
(b):
none of the discrete conformal structures (7), (8) and (9) can exist alone on an ideally triangulated surface with boundary, and the discrete conformal structures (7), (8) and (9) can not exist simultaneously in pairs on the same ideally triangulated surface with boundary.
(c):
the only mixed types of discrete conformal structures on an ideally triangulated surface with boundary are (4) (7), (5) (8) and (6) (9).
@@ Acknowledgements
This work was once presented on the workshop “Discrete Conformal Geometry” taking place at Wuhan University in China from June 21st to June 24th, 2024. During the workshop, Professor Feng Luo at Rutgers University and Professor Tian Yang at Texas A&M University suggested the relationships between the twisted generalized triangles and the discrete conformal structures on surfaces with boundary. The authors thank Professor Feng Luo and Professor Tian Yang for their suggestions. The research of the authors is supported by National Natural Science Foundation of China under grant no. 12471057.
@@ 2.1. Lorentzian cross product
For any two vectors $x,y\in\mathbb{H}^{2}$, there is a unique geodesic through $x$ and $y$. The arc length of the geodesic between $x$ and $y$ is defined to be the hyperbolic distance between $x$ and $y$, denoted by $d_{\mathbb{H}}(x,y)$. We use $\mathrm{Span}(x)$ and $\mathrm{Span}(x,y)$ to denote the vector subspace spanned $x$ and $x,y$, respectively. If $x\in\mathbb{R}^{3}$ is a space-like vector, then the Lorentzian complement $x^{\bot}:=\{w\in\mathbb{R}^{3}|x*w=0\}$ of $\mathrm{Span}(x)$ is a 2-dimension vector subspace of $\mathbb{R}^{3}$ and $x^{\bot}\cap\mathbb{H}^{2}$ is non-empty. This implies that $x^{\bot}\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$. And the hyperbolic distance between $x$ and $y\in\mathbb{H}^{2}$ is defined to be $d_{\mathbb{H}}(y,x^{\bot})=\inf\{d_{\mathbb{H}}(y,z)|z\in x^{\bot}\cap\mathbb{H}^{2}\}$. Let $x,y\in\mathbb{R}^{3}$ be space-like vectors. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is non-empty, then the hyperbolic distance between $x$ and $y$ is defined by $d_{\mathbb{H}}(x^{\bot},y^{\bot})=\inf\{d_{\mathbb{H}}(z,w)|z\in x^{\bot}\cap\mathbb{H}^{2},w\in y^{\bot}\cap\mathbb{H}^{2}\}$. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is empty, then $x^{\bot}$ and $y^{\bot}$ intersect in an angle $\angle(x^{\bot},y^{\bot})$ in $\mathbb{H}^{2}$.
@@ 2.3. Conformal variations of a right-angled hyperbolic hexagon
For a right-angled hyperbolic hexagon $\{ijk\}\in F$, there exists a unique hyper-ideal hyperbolic triangle $\{ijk\}$ corresponding to it with the vertices corresponding to the boundary arcs. We consider the effect of perturbing $f_{k}$ conformally while keeping $f_{i}$ and $f_{j}$ fixed, i.e., $\delta f_{i}=\delta f_{j}=0$. This is called an $f_{k}-$conformal variation by Glickenstein-Thomas [6]. In [6], Glickenstein-Thomas added a condition to a hyperbolic triangle in a closed surface, requiring that the vertex $v_{k}$ and face center $c_{ijk}$ of a hyperbolic triangle should be in a line including the vertex $\overline{v}_{k}$ of the new hyperbolic triangle under the $f_{k}-$conformal variation. We change the condition “in a line” in Glickenstein-Thomas [6] to the condition “in a 2-dimensional vector subspace of $\mathbb{R}^{3}$” for a hyper-ideal hyperbolic triangle in an ideally triangulated surface with boundary. In the Klein model, this is almost equivalent to the condition introduced by Glickenstein-Thomas, except that the vertices $v_{k}$ and $\overline{v}_{k}$ are spacelike. Please refer to Figure 2. Note that if the vertex $v_{i}$ of a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ is fixed, then the position of boundary component $i=v_{i}^{\bot}\cap\mathbb{H}^{2}$ will not move, while the arc length may change.
@@
(32) $\log\frac{\sinh^{2}d_{ij}}{\sinh^{2}d_{ji}}+\log\frac{\sinh^{2}d_{jk}}{\sinh^{2}d_{kj}}+\log\frac{\sinh^{2}d_{ki}}{\sinh^{2}d_{ik}}=0.$
Differentiating (32) with respect to $f_{i}$ gives $\frac{\partial}{\partial f_{i}}(\log\frac{\sinh^{2}d_{ki}}{\sinh^{2}d_{ik}})=-2$. Combining with (29) gives $\frac{\partial}{\partial f_{k}}(\log\frac{\sinh^{2}d_{ki}}{\sinh^{2}d_{ik}})=2$. Similarly, $\frac{\partial}{\partial f_{j}}(\log\frac{\sinh^{2}d_{jk}}{\sinh^{2}d_{kj}})=2$ and $\frac{\partial}{\partial f_{k}}(\log\frac{\sinh^{2}d_{jk}}{\sinh^{2}d_{kj}})=-2$. Then $c_{ij}+c_{jk}+c_{ki}=0$ by (30). Hence, $C_{ij}+C_{jk}+C_{ki}=0$ and $C_{rs}+C_{sr}=0$ for $\{r,s\}\subseteq\{i,j,k\}$. This gives the discrete conformal structure (6).
@@
This gives the discrete conformal structures (4) and (5).
@@
This gives the discrete conformal structures (7) and (8).
@@
Now we prove the claims (a), (b) and (c) in Theorem 1.4.
(1):
By the arguments above, the discrete conformal structures (4), (5) and (6) exist on $(\Sigma,\mathcal{T})$ only if $\frac{\sinh d_{ij}}{\sinh d_{ji}}>0$ for the edge $\{ij\}$. Thus it is possible that there exist only the discrete conformal structures (4), (5) or (6) on an ideally triangulated surface with boundary. However, the discrete conformal structures (4), (5) and (6) can not exist simultaneously in pairs on the same ideally triangulated surface with boundary.
Firstly, the different signs of $1+\alpha_{r}e^{2f_{r}}$ ensure that (4) and (5) can not exist simultaneously on the same ideally triangulated surface with boundary. Otherwise, due to the connectedness of $(\Sigma,\mathcal{T})$, there exists a right-angled hyperbolic hexagon $\{ijk\}\in F$ such that $l_{ij}$ is given by (4) and $l_{jk}$ is given by (5). Then $01+\alpha_{j}e^{2f_{j}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0. This is a contradiction.
Secondly, suppose that there exists a right-angled hyperbolic hexagon $\{ijk\}$ with $l_{ij}$ given by (6). Without loss of generality, we can assume $\frac{\partial H}{\partial f_{i}}\equiv 2,\ \frac{\partial H}{\partial f_{j}}\equiv-2$, otherwise change the orientation of the edge $\{ij\}$. By the arguments below (32), we see that $l_{jk},l_{ik}$ are also given by (6). Similar arguments also apply to the discrete conformal structures (4) and (5). Hence, (4) and (6), or (5) and (6) can not exist simultaneously on the same ideally triangulated surface with boundary. This gives the proof of the claim (a).
(2):
By the arguments above again, the discrete conformal structures (7), (8) and (9) exist in $(\Sigma,\mathcal{T})$ only if $\frac{\sinh d_{ij}}{\sinh d_{ji}}<0$ for the edge $\{ij\}$. This implies that the discrete conformal structures (7), (8) and (9) can not exist alone on an ideally triangulated surface with boundary. Otherwise, we have $\frac{\sinh d_{ij}}{\sinh d_{ji}}<0$$\frac{\sinh d_{ki}}{\sinh d_{ik}}<0$ and $\frac{\sinh d_{jk}}{\sinh d_{kj}}<0$ for a right-angled hyperbolic hexagon $\{ijk\}$, which implies $\frac{\sinh d_{ij}\sinh d_{jk}\sinh d_{ki}}{\sinh d_{ji}\sinh d_{ik}\sinh d_{kj}}<0$. This contradicts the condition (1). Moreover, the discrete conformal structures (7), (8) and (9) can not exist simultaneously in pairs on the same ideally triangulated surface with boundary. The proof is the same as that for the second part of claim (a). This gives the proof of the claim (b).
(3):
By the condition (1), if one of $\frac{\sinh d_{ij}}{\sinh d_{ji}}$$\frac{\sinh d_{ki}}{\sinh d_{ik}}$, $\frac{\sinh d_{jk}}{\sinh d_{kj}}$ is negative, then there is another unique one of $\frac{\sinh d_{ij}}{\sinh d_{ji}}$$\frac{\sinh d_{ki}}{\sinh d_{ik}}$, $\frac{\sinh d_{jk}}{\sinh d_{kj}}$ is also negative. For example, $\frac{\sinh d_{ij}}{\sinh d_{ji}}<0$, $\frac{\sinh d_{ki}}{\sinh d_{ik}}<0$ and $\frac{\sinh d_{jk}}{\sinh d_{kj}}>0$. This implies that two of the edge lengths $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}$ are induced by (7), (8) or (9), and the last one is induced by (4), (5) or (6). By the claim (b), the two of $l_{ij},l_{jk},l_{ki}$ with negative $\frac{\sinh d_{st}}{\sinh d_{ts}}$ are induced by the same one of (7), (8) and (9). By permutations and combinations, there are nine mixed types of discrete conformal structures, i.e., the discrete conformal structures (4) (8), (4) (9), (5) (7), (5) (9), (6) (7), (6) (8), (4) (7), (5) (8) and (6) (9). The first six mixed types of discrete conformal structures can not exist on the ideally triangulated surface with boundary $(\Sigma,\mathcal{T})$, while the last three can. The proof is also the same as that for the second part of claim (a). This proves the claim (c).
@@ 3. Relationships with 3-dimensional hyperbolic geometry
Denote the hyperbolic line dual to $O,v_{i},v_{j}$ by $L_{O},L_{i},L_{j}$ respectively. According to Theorem 1.4 (a), the discrete conformal structures (4), (5) and (6) can exist alone on an ideally triangulated surface with boundary. Hence, the three lateral generalized triangles with the cosine law from (4), (5) or (6) of a hyperbolic tetrahedron are the same.
@@
Now we give the geometric explanations of the discrete conformal structures (7), (8) and (9). According to Theorem 1.4 (b), the discrete conformal structures (7), (8) and (9) can not exist alone on an ideally triangulated surface with boundary. And we consider the mixed type of discrete conformal structures, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (4) and (7), (5) and (8) or (6) and (9). To ensure the condition (1) holds, there must be one of $l_{ij},l_{jk},l_{ki}$ from (4) and the other two from (7), one from (5) and the other two from (8), or one from (6) and the other two from (9). In other words, in a hyperbolic tetrahedron, there exist one lateral generalized triangle with the cosine law from (4) and the other two from (7), one lateral generalized triangle with the cosine law from (5) and the other two from (8), or one lateral generalized triangle with the cosine law from (6) and the other two from (9). For the discrete conformal structures (4), (5) and (6), the lateral generalized triangles are still shown in Figure 4. However, for the discrete conformal structures (7), (8) and (9), the lateral generalized triangles are some twisted hyperbolic triangles introduced by Roger-Yang [11]. Please refer to Figure 5. One can also refer to [1, 11] for more information on twisted hyperbolic triangles.
@@
Figure 6. Schematic diagram of four generalized hyperbolic tetrahedra and the corresponding lateral generalized triangles in the Kelin model.